# MatClaw: An Autonomous Code-First LLM Agent for End-to-End Materials Exploration

### 03 Apr 2026 v1

@@ 1 Introduction
Several materials-science agents have been proposed recently, including systems for autonomous electrochemistry (Zheng et al., 2025), specific simulation tasks (Liu et al., 2025), topological materials discovery (Zhang et al., 2025a), and first-principles VASP simulations (Xia et al., 2025). In adjacent fields, Coscientist (Boiko et al., 2023) and ChemCrow (Bran et al., 2024) demonstrated autonomous chemical research, El Agente (Zou et al., 2025) targeted quantum chemistry, and multi-agent frameworks for atomistic simulations have appeared (Vriza et al., 2026; Ansari and Moosavi, 2024).
@@
Pipeline-bounded architectures. Most agents are confined to a fixed set of supported software and predefined task sequences. For example, VASPilot (Liu et al., 2025) automates only VASP workflows, and Xia et al. (Xia et al., 2025) benchmark their agent on a fixed set of first-principles tasks. In practice, materials research routinely orchestrates multiple codes in a single workflow (e.g., DFT labeling with VASP, force field training with DeePMD-kit, and molecular dynamics simulations), and a pipeline-bounded agent cannot adapt to such multi-code workflows without substantial re-engineering.
@@ 3 Demonstration tasks and failure analysis
We validate MatClaw on a research project involving molecular dynamics (MD) simulations of monolayer (CIPS). CIPS is a van der Waals layered ferroelectric in which Cu atoms occupy a double-well potential along the out-of-plane direction; an applied electric field can drive Cu atoms from one potential well to the other, switching the local polarization. He et al. (He and others, 2023) trained a DeePMD potential for CIPS using the DP-GEN concurrent learning scheme (Zhang and others, 2020) and used it to simulate polarization switching under electric fields at elevated temperatures, observing spatially uncorrelated switching in which each Cu site nucleates independently—consistent with the nucleation-limited switching (NLS) picture. In conventional ferroelectrics, however, polarization reversal can also proceed through field-driven domain wall motion: walls behave as pinned elastic interfaces that advance by thermally activated creep at low drive and cross over to depinning at higher fields (Paruch and Guyonnet, 2013; Liu et al., 2016). This raises the question of whether CIPS, despite the NLS-like behavior reported at elevated temperatures, can access a more correlated wall-propagation regime under different conditions—specifically, whether lowering temperature to suppress stochastic thermal nucleation while increasing the electric field can drive a crossover from independent local flipping to sequential domain wall propagation. The boundary between these regimes in electric field–temperature parameter space has not been mapped.
@@
All three tasks use the pre-trained DeePMD model from He et al. (He and others, 2023) as the ground-truth potential rather than DFT calculations, reducing labeling cost from hours per frame (VASP) to seconds (teacher model inference). This substitution does not affect the test of agent capability: the workflow orchestration—supercell construction, MD submission, data conversion, active learning iteration, and convergence checking—is identical regardless of whether labels come from DFT or a teacher model. Tasks 2 and 3 use the teacher model directly for production MD.
@@ 3.1 Task 1: ML force field distillation via active learning
Training a machine-learning interatomic potential via active learning is a multi-stage iterative workflow: train an ensemble of models, explore configuration space with MD, identify high-uncertainty structures through inter-model disagreement, label them with a reference method, and retrain (Zhang and others, 2020). The central challenge is not the workflow orchestration itself, but the design of adequate sampling strategies—choosing simulation conditions (temperatures, pressures, simulation lengths) that cover the relevant physics, such as phase transitions and barrier crossings, rather than merely sampling near-equilibrium harmonic oscillations.
@@
To address this, we modified the task description in two ways. First, we prepended a “Background reading” section that provides the He et al. paper (He and others, 2023) and instructs the agent to extract the active learning methodology and persist key points as experience notes before proceeding with the task. Second, we added a single constraint: “Initial teacher MDs: at least 20 ps per temperature.” The rest of the prompt—workflow structure, stopping condition, output format—remained identical to the first attempt. The prepended section reads:
@@
Box 1. Experience note extracted by the agent from He et al. (He and others, 2023). Key parameters from He et al. for CuInP2S6 DP model training:
@@ 4 RAG benchmarks
We evaluate RAG effectiveness using three multiple-choice QA benchmark suites spanning a spectrum of library popularity (Table 3). The pymatgen suite (300 questions) is drawn from the MatTools benchmark (Liu and others, 2025). The VASP wiki suite (500 questions) and jobflow-remote suite (300 questions) were generated for this work using a similar methodology: an LLM reads each source or documentation file and produces structured multiple-choice questions with four options via a Pydantic schema, which are then validated and curated into balanced category sets. The VASP wiki questions cover four categories of domain knowledge (125 each): tag identification (which INCAR tag to use for a given task), physics recipe (which combination of settings achieves a given objective), error/restart handling (how to fix a specific error or restart a calculation), and tag dependency (which companion tags or constraints are required). The jobflow-remote questions cover four categories of API knowledge (75 each): function identification, parameter knowledge, return values, and usage patterns. Both source code QA and documentation QA are included because the agent needs both: correct Python API calls for workflow orchestration and correct simulation parameters for physically meaningful calculations.
@@
| Pymatgen code QA | Source code | 300 | MatTools (Liu and others, 2025) | — |

### 16 Apr 2026 v2

@@ 1 Introduction
Several materials-science agents have been proposed recently, including systems for autonomous electrochemistry (Zheng et al., 2025), specific simulation tasks (Liu et al., 2025a), topological materials discovery (Zhang et al., 2025a), and first-principles VASP simulations (Xia et al., 2025). In adjacent fields, Coscientist (Boiko et al., 2023) and ChemCrow (Bran et al., 2024) demonstrated autonomous chemical research, El Agente (Zou et al., 2025) targeted quantum chemistry, and multi-agent frameworks for atomistic simulations have appeared (Vriza et al., 2026; Ansari and Moosavi, 2024).
@@
Pipeline-bounded architectures. Most agents are confined to a fixed set of supported software and predefined task sequences. For example, VASPilot (Liu et al., 2025a) automates only VASP workflows, and Xia et al. (Xia et al., 2025) benchmark their agent on a fixed set of first-principles tasks. In practice, materials research routinely orchestrates multiple codes in a single workflow (e.g., DFT labeling with VASP, force field training with DeePMD-kit, and molecular dynamics simulations), and a pipeline-bounded agent cannot adapt to such multi-code workflows without substantial re-engineering.
@@ 3 Demonstration tasks and failure analysis
We validate MatClaw on a research project involving molecular dynamics (MD) simulations of monolayer (CIPS). CIPS is a van der Waals layered ferroelectric in which Cu atoms occupy a double-well potential along the out-of-plane direction; an applied electric field can drive Cu atoms from one potential well to the other, switching the local polarization. He et al. (He et al., 2023) trained a DeePMD potential for CIPS using the DP-GEN concurrent learning scheme (Zhang et al., 2020) and used it to simulate polarization switching under electric fields at elevated temperatures, observing spatially uncorrelated switching in which each Cu site nucleates independently—consistent with the nucleation-limited switching (NLS) picture. In conventional ferroelectrics, however, polarization reversal can also proceed through field-driven domain wall motion: walls behave as pinned elastic interfaces that advance by thermally activated creep at low drive and cross over to depinning at higher fields (Paruch and Guyonnet, 2013; Liu et al., 2016). This raises the question of whether CIPS, despite the NLS-like behavior reported at elevated temperatures, can access a more correlated wall-propagation regime under different conditions—specifically, whether lowering temperature to suppress stochastic thermal nucleation while increasing the electric field can drive a crossover from independent local flipping to sequential domain wall propagation. The boundary between these regimes in electric field–temperature parameter space has not been mapped.
@@
All three tasks use the pre-trained DeePMD model from He et al. (He et al., 2023) as the ground-truth potential rather than DFT calculations, reducing labeling cost from hours per frame (VASP) to seconds (teacher model inference). This substitution does not affect the test of agent capability: the workflow orchestration—supercell construction, MD submission, data conversion, active learning iteration, and convergence checking—is identical regardless of whether labels come from DFT or a teacher model. Tasks 2 and 3 use the teacher model directly for production MD.
@@ 3.1 Task 1: ML force field distillation via active learning
Training a machine-learning interatomic potential via active learning is a multi-stage iterative workflow: train an ensemble of models, explore configuration space with MD, identify high-uncertainty structures through inter-model disagreement, label them with a reference method, and retrain (Zhang et al., 2020). The central challenge is not the workflow orchestration itself, but the design of adequate sampling strategies—choosing simulation conditions (temperatures, pressures, simulation lengths) that cover the relevant physics, such as phase transitions and barrier crossings, rather than merely sampling near-equilibrium harmonic oscillations.
@@
To address this, we modified the task description in two ways. First, we prepended a “Background reading” section that provides the He et al. paper (He et al., 2023) and instructs the agent to extract the active learning methodology and persist key points as experience notes before proceeding with the task. Second, we added a single constraint: “Initial teacher MDs: at least 20 ps per temperature.” The rest of the prompt—workflow structure, stopping condition, output format—remained identical to the first attempt. The prepended section reads:
@@
Box 1. Experience note extracted by the agent from He et al. (He et al., 2023). Key parameters from He et al. for CuInP2S6 DP model training:
@@ 3.2 Task 2: Curie temperature prediction
On the predicted value.
The agent’s estimate K differs from the experimental bulk value of 315 K. The gap accumulates from three largely independent sources. First, the pretrained DeePMD potential used here He et al. (2023) itself predicts K for bulk , so the reference point against which the agent’s simulation should be compared is the potential’s own transition, not experiment. Second, the simulation is performed on a monolayer rather than bulk, removing interlayer van der Waals coupling and shifting downward. Third, the agent selected a 661 supercell (360 atoms), which is small enough that finite-size effects further reduce the apparent ; nonzero Cu polarization persists into the nominally paraelectric phase, confirming the signature of these effects. None of these affect the methodological point, but they illustrate that the absolute depends on setup choices (force-field provenance, dimensionality, supercell size) that the agent does not yet reason about and that still require expert audit.
@@ 4 RAG benchmarks
We evaluate RAG effectiveness using three multiple-choice QA benchmark suites spanning a spectrum of library popularity (Table 3). The pymatgen suite (300 questions) is drawn from the MatTools benchmark (Liu et al., 2025b). The VASP wiki suite (500 questions) and jobflow-remote suite (300 questions) were generated for this work using a similar methodology: an LLM reads each source or documentation file and produces structured multiple-choice questions with four options via a Pydantic schema, which are then validated and curated into balanced category sets. The VASP wiki questions cover four categories of domain knowledge (125 each): tag identification (which INCAR tag to use for a given task), physics recipe (which combination of settings achieves a given objective), error/restart handling (how to fix a specific error or restart a calculation), and tag dependency (which companion tags or constraints are required). The jobflow-remote questions cover four categories of API knowledge (75 each): function identification, parameter knowledge, return values, and usage patterns. Both source code QA and documentation QA are included because the agent needs both: correct Python API calls for workflow orchestration and correct simulation parameters for physically meaningful calculations.
@@
| Pymatgen code QA | Source code | 300 | MatTools (Liu et al., 2025b) | — |
@@ Acknowledgements
We thank Rahul Rao for pointing out the discrepancy between the agent’s predicted Curie temperature for and the experimental bulk value, which prompted the clarifying discussion in Section 3.2.